y'' + 2y' + y = x + e^(-x) , y(0)=0 , y'(0)=0

Solve the following 2nd ODE and find y(x)
Answer #1

d Yu + 2 y + y = (x+2) – Let D = 1 ► Dey + 20y +y =(sten) (D2+20+1 ) y = (x te*) for complimentary function D2+20+1=0 @ +1 )Colett a vecern) 1 x=(1+orex =/1-20 + 30²t.... ) x -= x - 2 T = x - - otu? using former og - 1 =D P1 = (2-2) + x2 * { By eq ④Y%) to Initial condition 966) 50 and Putting sto ond y zo in ea o= (0-2) + ē (9) dipte mentity en ® y = 1+ e* (2+x ) +(4+Get